# DeepBooTS: Dual‑Stream Residual Boosting for Drift‑Resilient Time‑Series Forecasting

### v1

@@ A.3 Concept Drift in TS forecasting
Classical methods include using sliding windows of recent data to regularly update models, thereby “forgetting” outdated patterns. Traditional statistical forecasters like ARIMA (piccolo1990distance) can be incrementally retrained on rolling windows, e.g., an incremental ARIMA scheme fits a new model whenever the current model’s validation error exceeds a threshold, signaling a drift. Another staple is change-point detection, where statistical tests (e.g. the Page-Hinkley test) monitor the data or error distribution and flag abrupt shifts once differences exceed a predefined threshold. For instance, the Error Intersection Approach (EIA) (baier2020handling) employs two models, one “stable” (more complex) and one “reactive” (simpler), and compares their performance. Furthermore, the Drift Detection Method (DDM) (korycki2023adversarial) monitors the classifier’s error rate distribution and raises an alarm when the error’s increase is statistically significant. These classical and statistical methods form the foundation of drift handling in forecasting, using either explicit retraining policies or specialized detectors to maintain model validity over time.
Online learning strategies embed concept drift adaptation within the forecasting models themselves. One example is FSNet (pham2023learning), which leverages complementary learning systems theory to pair a slow-learning base forecaster with fast-adapting components. Another line of work is OneNet (wen2023onenet), an online ensembling approach that dynamically combines two neural models: one specializes in capturing temporal dependencies within each series, and the other focuses on cross-series (covariate) relationships. Each of these deep learning techniques illustrates how integrating drift-awareness (through dual-model architectures, ensembling, or proactive adjustment) can improve TS forecasting performance in online.
@@ B.1 Proof of Equation 1
Equation 21 demonstrates a triangular relationship between the loss, bias, and variance of the model when the noise term is not taken into account, as shown in Fig. 3. For modern complex deep learning models, their biases are typically very low (goodfellow2016deep; zhang2021understanding). Here, we focus on how to reduce their variance to alleviate concept drift. In Equation 9, the final model is a weighted sum of $L$ simple models, with the weights determined by the performance of the previous model.
@@ C.1 Commonly used TS datasets
The information of the experiment datasets used in this paper are summarized as follows: (1) Electricity Transformer Temperature (ETT) dataset (Zhou2021Informer), which contains the data collected from two electricity transformers in two separated counties in China, including the load and the oil temperature recorded every 15 minutes (ETTm) or 1 hour (ETTh) between July 2016 and July 2018. (2) Electricity (ECL) dataset  [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 collects the hourly electricity consumption of 321 clients (each column) from 2012 to 2014. (3) Exchange (lai2018modeling) records the current exchange of 8 different countries from 1990 to 2016. (4) Traffic dataset  [^2] http://pems.dot.ca.gov records the occupation rate of freeway system across State of California measured by 861 sensors. (5) Weather dataset  [^3] https://www.bgc-jena.mpg.de/wetter records every 10 minutes for 21 meteorological indicators in Germany throughout 2020. (6) Solar-Energy (lai2018modeling) documents the solar power generation of 137 photovoltaic (PV) facilities in the year 2006, with data collected at 10-minute intervals. (7) The PEMS dataset (liu2022scinet) comprises publicly available traffic network data from California, collected within 5-minute intervals and encompassing 358 attributes. (8) Illness (ILI) dataset  [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html describes the influenza-like illness patients in the United States between 2002 and 2021, which records the ratio of patients seen with illness and the total number of the patients. The detailed statistics information of the datasets is shown in Table 4.
@@ C.2 Monash TS Forecasting Datasets
(1) Saugeenday dataset  [^5] https://zenodo.org/records/4656058 contains a single very long time series representing the daily mean flow of the Saugeen River at Walkerton in cubic meters per second and the length of this time series is 23741. (2) Sunspot dataset  [^6] https://www.kaggle.com/datasets/robervalt/sunspots contains monthly numbers of sunspots, as from the World Data Center, aka SIDC, between 1749 and 2018 with a total observation of 3240 months. (3) M4 dataset (MAKRIDAKIS202054) is a collection of 100,000 time series used in the fourth Makridakis Prediction Contest. The dataset consists of a time series of annual, quarterly, monthly, and other weekly, daily, and hourly data. In this paper, we utilize the hourly version of the M4 dataset and standardize its length to 768. (4) NN5 dataset  [^7] http://www.neural-forecasting-competition.com/downloads/NN5/datasets/download.htm contains daily time series originated from the observation of daily withdrawals at 111 randomly selected different cash machines at different locations within England. (5) Oikolab Weather dataset  [^8] https://docs.oikolab.com contains hundreds of terabytes of weather data, all of them are post-processed data from national weather agencies. (6) US Births dataset  [^9] https://cran.r-project.org/web/packages/mosaicData provides birth rates and related data across the 50 states and DC from 2016 to 2021, with a total observation of 7305. (7) Rideshare dataset (godahewa2021monash) comprises diverse hourly time series representations of attributes pertinent to Uber and Lyft rideshare services across multiple locations in New York. The data spans from November 26th, 2018, to December 18th, 2018. All the above data sets are collected in the Monash library [^10] https://forecastingdata.org.
@@ C.3 Large-Scale Time Series Datasets
(1) Milano (barlacchi2015multi) is provided by SKIL [^11] http://jol.telecomitalia.com/jolskil of Telecom Italia. The dataset was collected from November 1, 2013, to January 1, 2014, and the data is aggregated into 10-minute intervals over the whole city of Milan (62 days, 300 million records, about 19 GB). Milano records the time of the interaction and the specific radio base station that managed it. The dataset contains 10,000 nodes, and the length of each node is 1498. The ratio of training, validation and test datasets is 10:2:3.
@@
(3) C2TM (Modeling15Chen) makes an analysis on week-long traffic generated by a large population of people in a median-size city of China. C2TM analyses make use of request-response records extracted from traffic at the city scale, consisting of individuals’ activities during a continuous week (actually eight days from Aug. 19 to Aug. 26, 2012), with accurate timestamp and location information indicated by connected cellular base stations. C2TM contains 13269 nodes, each with a length of 192. The ratio of training, validation and test datasets is 5:1:2.
@@ Appendix D Implementation Details
The model undergoes training utilizing the ADAM optimizer (kingma2014adam) and minimizing the Mean Squared Error (MSE) loss function. The training process is halted prematurely, typically within 10 epochs. The DeepBooTS architecture solely comprises the embedding layer and backbone architecture, devoid of any additional introduced hyperparameters. Refer to Appendix I for the hyperparameter sensitivity analysis. During model validation, two evaluation metrics are employed: MSE and Mean Absolute Error (MAE).
For the large-scale datasets, we initially employ the RSS algorithm in (liang2024act) to perform data sampling. In this algorithm, we random decompose the entire large graph into 17 and 25 sub-graphs for the CBS and Milano datasets, respectively. As in previous works (wang2024timexer; wang2024timemixer; liu2023itransformer; nie2022time; wu2021autoformer; Zhou2021Informer; liang2023does), the model is trained using the ADAM optimizer (kingma2014adam) under input-36-output-24 setting. The learning rate is set to 1e-4 with MSE loss. The batch size is set to 64 for training and 1 for testing, and the maximum number of epochs is set to 10.
@@ Appendix F Performance Gains on Large-Scale TS Datasets
We also evaluated the performance gain when DeepBooTS is configured with other models, including PatchTST (nie2022time), STID (shao2022spatial), GWNet (wu2019graph), Autoformer (wu2021autoformer) and Informer (Zhou2021Informer). As shown in Fig. 10, DeepBooTS brings significant performance gains to other models in large-scale TS datasets. DeepBooTS consistently lowers forecasting error across all five backbones and three datasets, with relative gains ranging from small (2% on C2TM) to massive (up to 94% on CBS), highlighting its ability to stabilize and sharpen predictions even for models struggling under heavy distribution shifts.
@@ Appendix I Hyperparameter Sensitivity
The full results for univariate TS forecasting are presented in Table 12. As other models, e.g., iTransformer (liu2023itransformer), PatchTST (nie2022time) and Crossformer (zhang2022crossformer) do not offer performance information for all prediction lengths, we compare our method with those that provide comprehensive performance analysis, including Periodformer (liang2023does), FEDformer (zhou2022fedformer), Autoformer (wu2021autoformer), Informer (Zhou2021Informer), LogTrans (li2019enhancing) and Reformer (Kitaev2020Reformer). Despite Periodformer being a model that determines optimal hyperparameters through search, the proposed method outperforms benchmark approaches by achieving the highest count of leading terms across various prediction lengths. This reaffirms the effectiveness of DeepBooTS.
@@
For clarity and comparison among different models, we present supplementary prediction showcases for three representative datasets in Fig. 15, 16, and 17. Visualization of different models for qualitative comparisons. Prediction cases from the Traffic, Electricity and Weather datasets. These showcases correspond to predictions made by the following models: iTransformer (liu2023itransformer) and PatchTST (nie2022time), DLinear (zeng2023transformers), Autoformer (wu2021autoformer) and Informer (Zhou2021Informer). Among the various models considered, the proposed DeepBooTS stands out for its ability to predict future series variations with exceptional precision, demonstrating superior performance.

### v2

@@ A.3 Concept Drift in TS forecasting
Classical methods include using sliding windows of recent data to regularly update models, thereby “forgetting” outdated patterns. Traditional statistical forecasters like ARIMA (Piccolo 1990) can be incrementally retrained on rolling windows, e.g., an incremental ARIMA scheme fits a new model whenever the current model’s validation error exceeds a threshold, signaling a drift. Another staple is change-point detection, where statistical tests (e.g. the Page-Hinkley test) monitor the data or error distribution and flag abrupt shifts once differences exceed a predefined threshold. For instance, the Error Intersection Approach (EIA) (Baier et al. 2020) employs two models, one “stable” (more complex) and one “reactive” (simpler), and compares their performance. Furthermore, the Drift Detection Method (DDM) (Korycki and Krawczyk 2023) monitors the classifier’s error rate distribution and raises an alarm when the error’s increase is statistically significant. These classical and statistical methods form the foundation of drift handling in forecasting, using either explicit retraining policies or specialized detectors to maintain model validity over time.
Online learning strategies embed concept drift adaptation within the forecasting models themselves. One example is FSNet (Pham et al. 2023), which leverages complementary learning systems theory to pair a slow-learning base forecaster with fast-adapting components. Another line of work is OneNet (Wen et al. 2023), an online ensembling approach that dynamically combines two neural models: one specializes in capturing temporal dependencies within each series, and the other focuses on cross-series (covariate) relationships. Each of these deep learning techniques illustrates how integrating drift-awareness (through dual-model architectures, ensembling, or proactive adjustment) can improve TS forecasting performance in online.
@@ B.1 Proof of Equation 1
Equation 21 demonstrates a triangular relationship between the loss, bias, and variance of the model when the noise term is not taken into account, as shown in Fig. 3. For modern complex deep learning models, their biases are typically very low (Goodfellow et al. 2016; Zhang et al. 2021). Here, we focus on how to reduce their variance to alleviate concept drift. In Equation 9, the final model is a weighted sum of $L$ simple models, with the weights determined by the performance of the previous model.
@@ C.1 Commonly used TS datasets
The information of the experiment datasets used in this paper are summarized as follows: (1) Electricity Transformer Temperature (ETT) dataset (Zhou et al. 2021), which contains the data collected from two electricity transformers in two separated counties in China, including the load and the oil temperature recorded every 15 minutes (ETTm) or 1 hour (ETTh) between July 2016 and July 2018. (2) Electricity (ECL) dataset  [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 collects the hourly electricity consumption of 321 clients (each column) from 2012 to 2014. (3) Exchange (Lai et al. 2018) records the current exchange of 8 different countries from 1990 to 2016. (4) Traffic dataset  [^2] http://pems.dot.ca.gov records the occupation rate of freeway system across State of California measured by 861 sensors. (5) Weather dataset  [^3] https://www.bgc-jena.mpg.de/wetter records every 10 minutes for 21 meteorological indicators in Germany throughout 2020. (6) Solar-Energy (Lai et al. 2018) documents the solar power generation of 137 photovoltaic (PV) facilities in the year 2006, with data collected at 10-minute intervals. (7) The PEMS dataset (Liu et al. 2022a) comprises publicly available traffic network data from California, collected within 5-minute intervals and encompassing 358 attributes. (8) Illness (ILI) dataset  [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html describes the influenza-like illness patients in the United States between 2002 and 2021, which records the ratio of patients seen with illness and the total number of the patients. The detailed statistics information of the datasets is shown in Table 4.
@@ C.2 Monash TS Forecasting Datasets
(1) Saugeenday dataset  [^5] https://zenodo.org/records/4656058 contains a single very long time series representing the daily mean flow of the Saugeen River at Walkerton in cubic meters per second and the length of this time series is 23741. (2) Sunspot dataset  [^6] https://www.kaggle.com/datasets/robervalt/sunspots contains monthly numbers of sunspots, as from the World Data Center, aka SIDC, between 1749 and 2018 with a total observation of 3240 months. (3) M4 dataset (Makridakis et al. 2020) is a collection of 100,000 time series used in the fourth Makridakis Prediction Contest. The dataset consists of a time series of annual, quarterly, monthly, and other weekly, daily, and hourly data. In this paper, we utilize the hourly version of the M4 dataset and standardize its length to 768. (4) NN5 dataset  [^7] http://www.neural-forecasting-competition.com/downloads/NN5/datasets/download.htm contains daily time series originated from the observation of daily withdrawals at 111 randomly selected different cash machines at different locations within England. (5) Oikolab Weather dataset  [^8] https://docs.oikolab.com contains hundreds of terabytes of weather data, all of them are post-processed data from national weather agencies. (6) US Births dataset  [^9] https://cran.r-project.org/web/packages/mosaicData provides birth rates and related data across the 50 states and DC from 2016 to 2021, with a total observation of 7305. (7) Rideshare dataset (Godahewa et al. 2021) comprises diverse hourly time series representations of attributes pertinent to Uber and Lyft rideshare services across multiple locations in New York. The data spans from November 26th, 2018, to December 18th, 2018. All the above data sets are collected in the Monash library [^10] https://forecastingdata.org.
@@ C.3 Large-Scale Time Series Datasets
(1) Milano (Barlacchi et al. 2015) is provided by SKIL [^11] http://jol.telecomitalia.com/jolskil of Telecom Italia. The dataset was collected from November 1, 2013, to January 1, 2014, and the data is aggregated into 10-minute intervals over the whole city of Milan (62 days, 300 million records, about 19 GB). Milano records the time of the interaction and the specific radio base station that managed it. The dataset contains 10,000 nodes, and the length of each node is 1498. The ratio of training, validation and test datasets is 10:2:3.
@@
(3) C2TM (Xiaming Chen et al. 2015) makes an analysis on week-long traffic generated by a large population of people in a median-size city of China. C2TM analyses make use of request-response records extracted from traffic at the city scale, consisting of individuals’ activities during a continuous week (actually eight days from Aug. 19 to Aug. 26, 2012), with accurate timestamp and location information indicated by connected cellular base stations. C2TM contains 13269 nodes, each with a length of 192. The ratio of training, validation and test datasets is 5:1:2.
@@ Appendix D Implementation Details
The model undergoes training utilizing the ADAM optimizer (Kingma and Ba 2015) and minimizing the Mean Squared Error (MSE) loss function. The training process is halted prematurely, typically within 10 epochs. The DeepBooTS architecture solely comprises the embedding layer and backbone architecture, devoid of any additional introduced hyperparameters. Refer to Appendix I for the hyperparameter sensitivity analysis. During model validation, two evaluation metrics are employed: MSE and Mean Absolute Error (MAE).
For the large-scale datasets, we initially employ the RSS algorithm in (Liang et al. 2024a) to perform data sampling. In this algorithm, we random decompose the entire large graph into 17 and 25 sub-graphs for the CBS and Milano datasets, respectively. As in previous works (Wang et al. 2024b, a; Liu et al. 2023; Nie et al. 2022; Wu et al. 2021; Zhou et al. 2021; Liang et al. 2023), the model is trained using the ADAM optimizer (Kingma and Ba 2015) under input-36-output-24 setting. The learning rate is set to 1e-4 with MSE loss. The batch size is set to 64 for training and 1 for testing, and the maximum number of epochs is set to 10.
@@ Appendix F Performance Gains on Large-Scale TS Datasets
We also evaluated the performance gain when DeepBooTS is configured with other models, including PatchTST (Nie et al. 2022), STID (Shao et al. 2022), GWNet (Wu et al. 2019), Autoformer (Wu et al. 2021) and Informer (Zhou et al. 2021). As shown in Fig. 10, DeepBooTS brings significant performance gains to other models in large-scale TS datasets. DeepBooTS consistently lowers forecasting error across all five backbones and three datasets, with relative gains ranging from small (2% on C2TM) to massive (up to 94% on CBS), highlighting its ability to stabilize and sharpen predictions even for models struggling under heavy distribution shifts.
@@ Appendix I Hyperparameter Sensitivity
The full results for univariate TS forecasting are presented in Table 12. As other models, e.g., iTransformer (Liu et al. 2023), PatchTST (Nie et al. 2022) and Crossformer (Zhang and Yan 2022) do not offer performance information for all prediction lengths, we compare our method with those that provide comprehensive performance analysis, including Periodformer (Liang et al. 2023), FEDformer (Zhou et al. 2022), Autoformer (Wu et al. 2021), Informer (Zhou et al. 2021), LogTrans (Li et al. 2019) and Reformer (Kitaev et al. 2020). Despite Periodformer being a model that determines optimal hyperparameters through search, the proposed method outperforms benchmark approaches by achieving the highest count of leading terms across various prediction lengths. This reaffirms the effectiveness of DeepBooTS.
@@
For clarity and comparison among different models, we present supplementary prediction showcases for three representative datasets in Fig. 15, 16, and 17. Visualization of different models for qualitative comparisons. Prediction cases from the Traffic, Electricity and Weather datasets. These showcases correspond to predictions made by the following models: iTransformer (Liu et al. 2023) and PatchTST (Nie et al. 2022), DLinear (Zeng et al. 2023), Autoformer (Wu et al. 2021) and Informer (Zhou et al. 2021). Among the various models considered, the proposed DeepBooTS stands out for its ability to predict future series variations with exceptional precision, demonstrating superior performance.